# MoD2T:Model-Data-Driven Motion-Static Object Tracking Method

### v1

@@ III-C Motion Judgment
Among them, $I_{1},I_{2}$ represent two background feature tensors, $\mu_{I_{1}},\mu_{I_{2}}$ represent the mean of the two tensors, $\sigma_{I_{1}}^{2},\sigma_{I_{2}}^{2}$ represent the variance of the two tensors, $\sigma_{I_{1}I_{2}}$ represents the covariance of the two tensors, $r_{1},r_{2}$ is a constant, used to avoid zero denominator, and $\beta_{d}$ is the maximum distance that the preset stationary object may move in the image. In this paper, $\beta_{d}$ is usually set based on experience to be 0.01 times the smaller one in the length and width of the image.
@@ IV MoD2T
In the previous sections, we discussed the advantages and limitations of both traditional methods and deep learning methods. The objective of this paper is to parallelize and integrate these two approaches to achieve improved results. We propose a model-data-driven motion-static object tracking method (MoD2T) that combines the outputs of traditional method and deep learning method to enhance tracking accuracy and robustness. The traditional method used in this paper proposes an improvement to the Gaussian model proposed in [2], and through experiments, it demonstrates that this method, when combined with the OC-SORT[46], achieves satisfactory results in many environments. For the specific implementation details, please refer to the appendix. Inspired by the work in [38], which demonstrated the benefits of weighted fusion for multiple tracking methods, we design a fusion scheme for the methods employed in this paper. Let’s denote the two tracking methods as $M=\left\{M^{tra},M^{deep}\right\}$. For each frame $I_{t},t=1,2,...,T$, where $T$ represents the total number of frames in the video, these methods generate a series of the bounding box (bbox) ($\widehat{B}_{t}^{i}$) and motion assessment values ($\widehat{\varTheta}_{t}^{i}$) , i.e. $i=1,...,n$, where $n$ represents the number of detections in the $t$ frame. The fusion method employed in this paper combines their outputs $\widehat{B}_{t}=(\widehat{B}_{t}^{tra},\widehat{B}_{t}^{deep})$ and $\widehat{\varTheta}_{t}=(\widehat{\varTheta}_{t}^{tra},\widehat{\varTheta}_{t}^{deep})$ to obtain the fused tracking outputs $\widehat{B}_{t}^{fusion},\widehat{\varTheta}_{t}^{fusion}$. Considering the performance issues associated with traditional methods, our fusion strategy may undergo certain adjustments based on the varying performance of the traditional methods. In a comprehensive overview, the procedural workflow of MoD2T is illustrated in Fig.4. Subsequent to the independent processing of image data using two distinct methodologies, we apply the motion classification technique outlined in section III to the outcomes derived from the deep learning method. The precise fusion strategy will be contingent upon the efficacy of the traditional approach.
@@ IV-A Traditional Methods Evaluation
Our evaluation is not based on the overall results of the entire image, but rather on evaluating individual tracking results. This is because traditional methods can produce erroneous results when performing morphological operations such as erosion and dilation, leading to two objects being mistakenly identified as one. Therefore, we focus on the connected components within the bounding boxes of each moving object obtained by the traditional method for evaluation. For each motion pixel $p$, we examine its surrounding neighboring pixels within a small range, typically set to three pixels, inspired by the concept of a 3x3 convolution kernel commonly used in Convolutional Neural Networks (CNN). This approach is akin to the local receptive field employed by CNN, designed to capture local information in images. If additional motion pixels are detected within this specified range, the current pixel is designated as a new connected component and is assigned a unique label. Subsequently, we analyze the resulting labeled image and quantify the number of connected components. This choice of a small range, specifically three pixels, draws inspiration from the effectiveness of 3x3 convolutional kernels in CNN, ensuring a balance between computational efficiency and the model’s ability to identify and label connected components of motion pixels. We sum the pixels in these connected components, and if the pixel value exceeds $S_{min}$, we consider it as a valid connected component. The specific expression of this calculation can be represented by the (10):
@@
When $n\geq 2,$ we consider that the traditional method has not worked properly. By calculating and observing the number of connected components obtained by the traditional method, we can assess its performance in handling situations involving close or overlapping objects. This evaluation method allows for a more detailed understanding of the method’s performance on individual tracking results and provides an assessment of its effectiveness. One more thing, in this paper, $S_{min}$ generally set to one-fifth of the bbox area. A typical example of the effectiveness of a traditional tracking method is illustrated in Fig. 3.
@@
Figure 4: The comprehensive architecture of MoD2T is presented, allowing for flexibility in the integration of various deep learning-based methods. Specifically, this study employs YOLOv8+StrongSORT and YOLOv8+DeepOCSORT as instances of such methods. Fig.7 illustrates the traditional method, while Fig.1 focuses on the motion judgment aspect. Within the algorithmic effectiveness assessment module depicted in the figure, the region enclosed by purple dashed lines signifies multiple potential outputs of this module. The notation $p_{t}$ within the diagram represents the pixels that traditional algorithms may identify as moving objects.
@@ IV-B Fusion Strategies in the Presence of Effective Traditional Methods
where, $\alpha_{k}^{b}$ is the confidence level, given by humans, and MOTA is a common evaluation indicator for multi-objective tracking methods[47]. The reason for choosing MOTA is that it mainly reflects whether the object has been tracked, and traditional methods often miss detection of the object. Therefore, using this indicator as the standard for calculating fusion parameters is reasonable. Since deep learning methods track both moving and stationary objects, the tracking results from deep learning often contain more detections compared to traditional methods. In cases where there are no corresponding detections from the traditional method for the deep learning results, we directly adopt the results from the deep learning method. In other words, we set $w_{deep}^{b}=1,w_{tra}^{b}=0$ in this cases. This ensures that the deep learning results are considered as the primary source of tracking information when there is no corresponding result from the traditional methods. The value $\varTheta$, which indicates whether an object is in motion or not, is also calculated using a weighted average in this paper, as shown in (13). The calculation of $\varTheta^{deep}$ is defined in (7). For $\varTheta_{i}^{tra}$, when there is no corresponding detection from the traditional method for the deep learning results, it is set to 1; otherwise, it is set to 0. Meanwhile, if the traditional method does not work properly, its value will be shown in equation (16)
@@
Subscript $i$ means the ID of different tracking objects. Similar to the weights for bbox, the weights for $\varTheta$ are also determined based on the performance of each method on the same dataset. The weight coefficients can be calculated as shown in (14),
@@
IV-C Object Motion Assessment: MVF1 and its Elements
Here, $MVTP$ signifies the count of samples accurately classified as either moving or stationary objects, $MVFP$ denotes the count of samples inaccurately labeled as moving objects when they are stationary or as stationary objects when they are in motion, and $MVFN$ represents the count of samples absent from the ground truth. $BF$ stands for the balance factor, a parameter adjustable to enhance adaptability. The introduction of $BF$ is motivated by the metric’s primary purpose of measuring the accuracy in determining the motion state of objects. Ideally, the impact of tracking or detection performance should be disregarded. Nevertheless, given that motion-state determination relies on tracking, this metric retains a certain influence from the detection process.
$BF=\frac{1}{2}\cdot(\text{precision}+\text{recall})$ (16)
IV-D Fusion Strategy in the Presence of Malfunctioning Traditional Methods
@@ V-A Setting
The KITTI dataset [50] has achieved wide adoption within the autonomous driving and computer vision research communities. It encompasses data collected from diverse sensors like cameras and LiDAR mounted on vehicles. The dataset presents an array of scenes annotated with object bounding boxes and motion trajectories for various entities, including vehicles and pedestrians, thereby rendering it a fitting choice for the evaluation of multi-object tracking methodologies. Moreover, our investigation extends to selected scenarios from the UAVDT [51] and MOT17 [13] datasets. The UAVDT dataset is meticulously tailored for video analysis of unmanned aerial vehicles, featuring captured video sequences across a spectrum of scenarios. On the other hand, the MOT17 dataset emphasizes multi-object tracking pertaining to pedestrians, offering video footage captured from different vantage points along with annotations detailing the motion trajectories of multiple objects. To underscore the versatility and efficacy of our method, we subject these UAVDT and MOT17 scenes to supplementary evaluation, bolstered by augmented annotations. These curated scenes introduce heightened complexity and intricate tracking scenarios, thereby affording a comprehensive assessment of the method’s resilience and adaptability. By subjecting our method to scrutiny across diverse datasets, we attain a holistic comprehension of its performance in multi-object tracking tasks and corroborate its efficacy across a plethora of scenarios. Our assessment strategy was systematic, entailing an exhaustive performance analysis conducted half of KITTI’s training set for training and estimating some parameters, and use the other half for testing. Given the need for supplemental annotations concerning object motion attributes within the MOT17 and UAVDT datasets, the entirety of these datasets was not embraced in this inquiry. Instead, our evaluation was concentrated on representative scenes, specifically MOT17 sequences (03, 07) and UAVDT sequences (M101, M202, M204, M402, M601, M703, M802). The outcomes of these exacting evaluations have been scrupulously collated in the forthcoming table. In a bid to furnish a comprehensive insight into the effectiveness and prowess of our proposed approach, we partitioned the KITTI dataset into two discrete segments for the purpose of testing. One segment encompasses sequences (0001,0017-0024), characterized by static or slowly moving cameras, while the other segment encapsulates scenarios involving swifter camera motions and accelerated object movements. One more thing, although using the test datasets for calculating the fusion weights, it is anticipated that this has minimal impact on the results. Nevertheless, in order to ensure the fairness of our experiments, this paper also employed additional datasets for a rough estimation of these weights. Specifically, for KITTI, the additional data encompassed the entire training set, which was further categorized based on the speed of motion. For MOT17, the selected datasets were 04 and 10, and for UAVDT, it included M203, M403, and M701. It is important to emphasize that although these datasets underwent additional annotation and can be utilized for computing tracking performance metrics and MVF1 scores, they were not used in the method’s performance testing in this study. Metrics. We use the metrics MOTA, IDF1,AssA and HOTA to evaluate tracking performance[47, 52, 53]. MOTA primarily measures the completeness and robustness of tracking, IDF1 focuses on the accuracy and precision of targets, while HOTA comprehensively considers multiple aspects of performance, including both target and trajectory-level accuracy. When comparing methods and evaluating performance, these metrics can assist researchers in gaining a comprehensive understanding of the strengths, weaknesses, and applicability of the methods. Additionally, we use the MVF1 to measure the accuracy of object motion state predictions. The calculation method for the MVF1 is referenced in (15). Implementation Details. We utilize OSNet[45] as the network for extracting background appearance features. For testing on the UAVDT and KITTI , we employ osnet pre-trained on VeRi[54], while for evaluating the MOT17 , we use weights pre-trained on Market1501[55]. Additionally, to conserve computational resources, the same osnet network is used in the deep learning-based MOT method adopted in this study. The criterion used to determine correctness is the Intersection over Union (IoU). A prediction is considered correct when the IoU between the method’s prediction and the ground truth is greater than 0.5. The training method of YOLOv8 in MOT17 is similar to [56]. The tracker used in traditional methods is OC-SORT[46].
V-B Reasonability analysis of MVF1 design
@@
As depicted in the Fig.5, the dynamic changes observed in MVF1, driven by varying levels of MOTA, offer valuable insights into the nuanced impact of BF parameter configurations on the model’s ability to assess the relative motion states of tracked objects with respect to the ground. Furthermore, it is noteworthy that distinct MOTA values were generated by randomly removing entire sequences with identical ID numbers. Examining the blue curve, representing MVF1 values under varying BF parameters, a clear positive correlation emerges with increasing MOTA levels. This trend emphasizes the utility of strategic BF parameter adjustments in enhancing the model’s capacity to accurately discern the motion states of tracked objects relative to the ground. Notably, the performance improvement is most pronounced at higher MOTA values, underscoring the synergistic relationship between MOTA and BF parameter optimization. Conversely, the green curve, denoting the scenario where the BF parameter is set to 1, suggests a nuanced trade-off. While initial MVF1 values exhibit promise, a subsequent decline in performance is observed as MOTA increases. This nuanced relationship underscores the importance of balancing BF parameter settings, as extreme values may compromise tracking accuracy. The orange curve, corresponding to a BF parameter value of 0, reveals intriguing stability in MVF1 values across varying MOTA levels. This finding indicates that, in specific contexts, the deliberate exclusion of background filtering may be a viable choice, particularly when minimizing false negatives is a critical consideration. In essence, the result demonstrates that judicious BF parameter settings significantly contribute to the enhanced evaluative capacity of MVF1, offering a more nuanced assessment of the model’s proficiency in discerning the motion states of tracked objects in relation to the ground. The observed trends highlight the need for a context-aware approach to BF parameter tuning, emphasizing the delicate balance required for optimal tracking performance.

### v2

@@ III-D Motion State Judgment
Among them,$I_{1}$ and $I_{2}$ denote two background feature tensors, with $\mu_{I_{1}}$ and $\mu_{I_{2}}$ representing their respective means, $\sigma_{I_{1}}^{2}$ and $\sigma_{I_{2}}^{2}$ denoting their variances, and $\sigma_{I_{1}I_{2}}$ indicating the covariance between them. Constants $r_{1}$ and $r_{2}$ are introduced to prevent a zero denominator, and $\beta_{d}$ represents the maximum distance a preset stationary object may move within the image.
In our approach, the value of $\beta_{d}$ is typically empirically set to 0.01 times the smaller dimension (length or width) of the image. This choice ensures a proportional scaling of $\beta_{d}$ and serves as a small adjustment factor based on the image’s dimensions. The rationale behind this selection is to strike a balance by considering image size in computations while maintaining a relatively small value to prevent the introduction of overly significant weights. The choice of using 0.01 as a multiplier is likely an empirical decision aimed at achieving a practical balance, ensuring algorithm effectiveness while minimizing computational demands.
@@ IV MoD2T
In the previous sections, we discussed the advantages and limitations of both traditional methods and deep learning methods. The objective of this paper is to parallelize and integrate these two approaches to achieve improved results. We propose a model-data-driven motion-static object tracking method (MoD2T) that combines the outputs of traditional method and deep learning method to enhance tracking accuracy and robustness.
The traditional method used in this paper proposes an improvement to the Gaussian model proposed in [2], and through experiments, it demonstrates that this method, when combined with the OC-SORT[51], achieves satisfactory results in many environments. For the specific implementation details, please refer to the appendix. Inspired by the work in [40], which demonstrated the benefits of weighted fusion for multiple tracking methods, we design a fusion scheme for the methods employed in this paper.
The flowchart depicted in Figure 3 outlines the MoD2T process, where two tracking methods, the traditional method represented as $M^{tra}$ and the deep learning-based method denoted as $M^{deep}$, run simultaneously. For each frame $I_{t}$, where $t=1,2,...,T$ and $T$ is the total number of frames in the video, both methods generate a series of bounding boxes ($\widehat{B}_{t}^{i}$) and corresponding motion assessment values ($\widehat{\varTheta}_{t}^{i}$), where $i=1,...,n$ and $n$ represents the number of detections in frame $t$. This paper’s fusion approach combines their individual outputs, resulting in $\widehat{B}_{t}=(\widehat{B}_{t}^{tra},\widehat{B}_{t}^{deep})$ and $\widehat{\varTheta}{t}=(\widehat{\varTheta}_{t}^{tra},\widehat{\varTheta}_{t}^{deep})$. The fused tracking outputs, denoted as $\widehat{B}_{t}^{fusion}$ and $\widehat{\varTheta}_{t}^{fusion}$, are then obtained through this fusion process.
Considering the performance issues associated with traditional methods, our fusion strategy may undergo certain adjustments based on the varying performance of the traditional methods.
Subsequent to the independent processing of image data using two distinct methodologies, we apply the motion classification technique outlined in section III to the outcomes derived from the deep learning method. The precise fusion strategy will be contingent upon the efficacy of the traditional approach.
@@ IV-A Traditional Methods Evaluation
Our evaluation is not based on the overall results of the entire image, but rather on evaluating individual tracking results. This is because traditional methods can produce erroneous results when performing morphological operations such as erosion and dilation, leading to two objects being mistakenly identified as one. Therefore, we focus on the connected components within the bounding boxes of each moving object obtained by the traditional method for evaluation.
For each motion pixel $p$, we examine its surrounding neighboring pixels within a small range, typically set to three pixels, inspired by the concept of a 3x3 convolution kernel commonly used in Convolutional Neural Networks (CNN). This approach is akin to the local receptive field employed by CNN, designed to capture local information in images. If additional motion pixels are detected within this specified range, the current pixel is designated as a new connected component and is assigned a unique label. Subsequently, we analyze the resulting labeled image and quantify the number of connected components. This choice of a small range, specifically three pixels, draws inspiration from the effectiveness of 3x3 convolutional kernels in CNN, ensuring a balance between computational efficiency and the model’s ability to identify and label connected components of motion pixels.
We sum the pixels in these connected components, and if the pixel value exceeds $S_{min}$, we consider it as a valid connected component. The specific expression of this calculation can be represented by the (10):
@@
When $n\geq 2,$ we consider that the traditional method has not worked properly. By calculating and observing the number of connected components obtained by the traditional method, we can assess its performance in handling situations involving close or overlapping objects. This evaluation method allows for a more detailed understanding of the method’s performance on individual tracking results and provides an assessment of its effectiveness. One more thing, in this paper, $S_{min}$ generally set to one-fifth of the bbox area. A typical example of the effectiveness of a traditional tracking method is illustrated in Figure. 4.
@@ IV-B Fusion Strategies in the Presence of Effective Traditional Methods
where, $\alpha_{k}^{b}$ is the confidence level, given by humans, and MOTA is a common evaluation indicator for multi-objective tracking methods[52]. The reason for choosing MOTA is that it mainly reflects whether the object has been tracked, and traditional methods often miss detection of the object. Therefore, using this indicator as the standard for calculating fusion parameters is reasonable.
Since deep learning methods track both moving and stationary objects, the tracking results from deep learning often contain more detections compared to traditional methods. In cases where there are no corresponding detections from the traditional method for the deep learning results, we directly adopt the results from the deep learning method. In other words, we set $w_{deep}^{b}=1,w_{tra}^{b}=0$ in this cases. This ensures that the deep learning results are considered as the primary source of tracking information when there is no corresponding result from the traditional methods.
The value $\varTheta$, which indicates whether an object is in motion or not, is also calculated using a weighted average in this paper, as shown in (13). The calculation of $\varTheta^{deep}$ is defined in (7). For $\varTheta_{i}^{tra}$, when there is no corresponding detection from the traditional method for the deep learning results, it is set to 1; otherwise, it is set to 0. Meanwhile, if the traditional method does not work properly, its value will be shown in equation (16)
@@
Subscript $i$ means the ID of different tracking objects.
Similar to the weights for bbox, the weights for $\varTheta$ are also determined based on the performance of each method on the same dataset. The weight coefficients can be calculated as shown in (14),
@@
IV-C Fusion Strategy in the Presence of Malfunctioning Traditional Methods
@@
IV-D Object Motion Metrics: Motion State Validation F1 Score
Here, $MVTP$ signifies the count of samples accurately classified as either moving or stationary objects, $MVFP$ denotes the count of samples inaccurately labeled as moving objects when they are stationary or as stationary objects when they are in motion, and $MVFN$ represents the count of samples absent from the ground truth. $BF$ stands for the balance factor, a parameter adjustable to enhance adaptability.
The introduction of $BF$ is motivated by the metric’s primary purpose of measuring the accuracy in determining the motion state of objects. Ideally, the impact of tracking or detection performance should be disregarded. Nevertheless, given that motion-state determination relies on tracking, this metric retains a certain influence from the detection process.
$BF=\frac{1}{2}\cdot(\text{precision}+\text{recall})$ (17)
@@ V-A Setting
The KITTI dataset [55] has achieved wide adoption within the autonomous driving and computer vision research communities. It encompasses data collected from diverse sensors like cameras and LiDAR mounted on vehicles. The dataset presents an array of scenes annotated with object bounding boxes and motion trajectories for various entities, including vehicles and pedestrians, thereby rendering it a fitting choice for the evaluation of multi-object tracking methodologies.
Moreover, our investigation extends to selected scenarios from the UAVDT [56] and MOT17 [15] datasets. The UAVDT dataset is meticulously tailored for video analysis of unmanned aerial vehicles, featuring captured video sequences across a spectrum of scenarios. On the other hand, the MOT17 dataset emphasizes multi-object tracking pertaining to pedestrians, offering video footage captured from different vantage points along with annotations detailing the motion trajectories of multiple objects.
To underscore the versatility and efficacy of our method, we subject these UAVDT and MOT17 scenes to supplementary evaluation, bolstered by augmented annotations. These curated scenes introduce heightened complexity and intricate tracking scenarios, thereby affording a comprehensive assessment of the method’s resilience and adaptability. By subjecting our method to scrutiny across diverse datasets, we attain a holistic comprehension of its performance in multi-object tracking tasks and corroborate its efficacy across a plethora of scenarios.
Our assessment strategy was systematic, entailing an exhaustive performance analysis conducted half of KITTI’s training set for training and estimating some parameters, and use the other half for testing. Given the need for supplemental annotations concerning object motion attributes within the MOT17 and UAVDT datasets, the entirety of these datasets was not embraced in this inquiry. Instead, our evaluation was concentrated on representative scenes, specifically MOT17 sequences (03, 07) and UAVDT sequences (M101, M202, M204, M402, M601, M703, M802).
The outcomes of these exacting evaluations have been scrupulously collated in the forthcoming table. In a bid to furnish a comprehensive insight into the effectiveness and prowess of our proposed approach, we partitioned the KITTI dataset into two discrete segments for the purpose of testing. One segment encompasses sequences (0001,0017-0024), characterized by static or slowly moving cameras, while the other segment encapsulates scenarios involving swifter camera motions and accelerated object movements.
One more thing, although using the test datasets for calculating the fusion weights, it is anticipated that this has minimal impact on the results. Nevertheless, in order to ensure the fairness of our experiments, this paper also employed additional datasets for a rough estimation of these weights. Specifically, for KITTI, the additional data encompassed the entire training set, which was further categorized based on the speed of motion. For MOT17, the selected datasets were 04 and 10, and for UAVDT, it included M203, M403, and M701. It is important to emphasize that although these datasets underwent additional annotation and can be utilized for computing tracking performance metrics and MVF1 scores, they were not used in the method’s performance testing in this study.
Metrics. We use the metrics MOTA, IDF1,AssA and HOTA to evaluate tracking performance[52, 57, 58]. MOTA primarily measures the completeness and robustness of tracking, IDF1 focuses on the accuracy and precision of targets, while HOTA comprehensively considers multiple aspects of performance, including both target and trajectory-level accuracy. When comparing methods and evaluating performance, these metrics can assist researchers in gaining a comprehensive understanding of the strengths, weaknesses, and applicability of the methods.
Additionally, we use the MVF1 to measure the accuracy of object motion state predictions. The calculation method for the MVF1 is referenced in (15).
Implementation Details. We utilize OSNet[50] as the network for extracting background appearance features. For testing on the UAVDT and KITTI , we employ osnet pre-trained on VeRi[59], while for evaluating the MOT17 , we use weights pre-trained on Market1501[60]. Additionally, to conserve computational resources, the same osnet network is used in the deep learning-based MOT method adopted in this study. The criterion used to determine correctness is the Intersection over Union (IoU). A prediction is considered correct when the IoU between the method’s prediction and the ground truth is greater than 0.5. The training method of YOLOv8 in MOT17 is similar to [61]. The tracker used in traditional methods is OC-SORT[51]. Specifically, for the KITTI dataset, this paper divides each sequence in its training set into two equally large parts, with the first part used for training and the second part used for testing.
V-B Reasonability analysis of MVF1
@@
As depicted in the Figure.5, the dynamic changes observed in MVF1, driven by varying levels of MOTA, offer valuable insights into the nuanced impact of BF parameter configurations on the model’s ability to assess the relative motion states of tracked objects with respect to the ground. Furthermore, it is noteworthy that distinct MOTA values were generated by randomly removing entire sequences with identical ID numbers.
Examining the blue curve, representing MVF1 values under varying BF parameters, a clear positive correlation emerges with increasing MOTA levels. This trend emphasizes the utility of strategic BF parameter adjustments in enhancing the model’s capacity to accurately discern the motion states of tracked objects relative to the ground. Notably, the performance improvement is most pronounced at higher MOTA values, underscoring the synergistic relationship between MOTA and BF parameter optimization.
Conversely, the green curve, denoting the scenario where the BF parameter is set to 1, suggests a nuanced trade-off. While initial MVF1 values exhibit promise, a subsequent decline in performance is observed as MOTA increases. This nuanced relationship underscores the importance of balancing BF parameter settings, as extreme values may compromise tracking accuracy.
The orange curve, corresponding to a BF parameter value of 0, reveals intriguing stability in MVF1 values across varying MOTA levels. This finding indicates that, in specific contexts, the deliberate exclusion of background filtering may be a viable choice, particularly when minimizing false negatives is a critical consideration.
In essence, the result demonstrates that judicious BF parameter settings significantly contribute to the enhanced evaluative capacity of MVF1, offering a more nuanced assessment of the model’s proficiency in discerning the motion states of tracked objects in relation to the ground. The observed trends highlight the need for a context-aware approach to BF parameter tuning, emphasizing the delicate balance required for optimal tracking performance.